# Two-Timescale Optimization Framework for Decentralized Linear-Quadratic Optimal Control

### v1

@@ 4 Solving Inner Layer Optimization Problem
$\mathcal{L}({\rm svec}(W);X)=$ $\langle\tilde{d}_{k},{\rm svec}(W)\rangle+\sigma_{1}\left\|\mathcal{A}D_{1}{\rm svec}(W)+\tilde{b}_{k}\right\|^{2}+\sigma_{2}\left\|D_{1}{\rm svec}(W)-\widetilde{v}_{k}\right\|^{2}$
@@
$\theta(X)=\min_{{\rm svec}(W)}\mathcal{L}({\rm svec}(W);X)=\langle X_{1}+\dots+X_{M},T_{2}(V\otimes V){\rm vec}(Q)\rangle+\min_{{\rm svec}(W)}\hat{\theta}({\rm svec}(W))$ (29)
@@
$\hat{\theta}({\rm svec}(W))=\langle\tilde{d}_{k},{\rm svec}(W)\rangle+\sigma_{1}\left\|\mathcal{A}D_{1}{\rm svec}(W)+\tilde{b}_{k}\right\|^{2}+\sigma_{2}\left\|D_{1}{\rm svec}(W)-\widetilde{v}_{k}\right\|^{2}.$
Let
$\nabla\hat{\theta}({\rm svec}(W))=\tilde{d}_{k}+2\sigma_{1}(\mathcal{A}D_{1})^{\top}(\mathcal{A}D_{1}{\rm svec}(W)+\tilde{b}_{k})+2\sigma_{2}D_{1}^{\top}(D_{1}{\rm svec}(W)-\tilde{v}_{k})=0.$
Alternatively, it may be derived that
$\left(2\sigma_{1}(\mathcal{A}D_{1})^{\top}(\mathcal{A}D_{1})+2\sigma_{2}D_{1}^{\top}D_{1}\right){\rm svec}(W)=-2\sigma_{1}(\mathcal{A}D_{1})^{\top}\tilde{b}_{k}+2\sigma_{2}D_{1}^{\top}\tilde{v}_{k}-\tilde{d}_{k}.$
Denoting
$M=2\sigma_{1}(\mathcal{A}D_{1})^{\top}(\mathcal{A}D_{1})+2\sigma_{2}D_{1}^{\top}D_{1},\leavevmode\nobreak\ q_{k}=-2\sigma_{1}(\mathcal{A}D_{1})^{\top}\tilde{b}_{k}+2\sigma_{2}D_{1}^{\top}\tilde{v}_{k},$
@@
Substituting (30) into (29), it could potentially inferred that
$\theta(X)$ $=\langle X_{1}+\dots+X_{M},T_{2}(V\otimes V){\rm vec}(Q)\rangle+\left\langle\tilde{d}_{k},M^{-1}(q_{k}-\tilde{d}_{k})\right\rangle$
$+\sigma_{1}\left\|\mathcal{A}D_{1}M^{-1}(q_{k}-\tilde{d}_{k})+\tilde{b_{k}}\right\|^{2}+\sigma_{2}\left\|D_{1}M^{-1}(q_{k}-\tilde{d}_{k})-\tilde{v}_{k}\right\|^{2}$
$=\langle X_{1}+\dots+X_{M},T_{2}(V\otimes V){\rm vec}(Q)\rangle+\tilde{d}_{k}^{\top}\left(\sigma_{1}(\mathcal{A}D_{1}M^{-1})^{\top}(\mathcal{A}D_{1}M^{-1})+\sigma_{2}M^{-2}-M^{-1}\right)\tilde{d}_{k}$
$+\left\langle\tilde{d}_{k},M^{-1}q_{k}+2\sigma_{1}(\mathcal{A}D_{1}M^{-1})^{\top}\xi_{k}^{1}+2\sigma_{2}(D_{1}M^{-1})^{\top}\xi_{k}^{2}\right\rangle+\|\xi_{k}^{1}\|^{2}+\|\xi_{k}^{2}\|^{2},$
@@
$\xi_{k}^{1}=-\mathcal{A}D_{1}M^{-1}q_{k}-\tilde{b}_{k},\leavevmode\nobreak\ \xi_{k}^{2}=\tilde{v}_{k}-D_{1}M^{-1}q_{k}.$
@@
$\max_{X=(X_{0},\dots,X_{M})}\leavevmode\nobreak$ $\langle X_{1}+\dots+X_{M},T_{2}(V\otimes V){\rm vec}(Q)\rangle+\left\langle\tilde{d}_{k},\Omega_{k}\right\rangle+\tilde{d}_{k}^{\top}\Psi\tilde{d}_{k}$ (31)
${\rm s.t.}\leavevmode\nobreak$ $X_{0}\in\widetilde{\Gamma}_{+}^{p},$
$X_{i}\in\widetilde{\Gamma}_{+}^{n},\leavevmode\nobreak\ i=1,\dots,M.$
@@
$\Theta(X_{0},X_{1},\dots,X_{M})$ $=-\langle X_{1}+\dots+X_{M},T_{2}(V\otimes V){\rm vec}(Q)\rangle-\left\langle\mathcal{L}_{k}(X_{0},\dots,X_{M}),\Omega_{k}\right\rangle$
$+\left\|\sqrt{-\Psi}\mathcal{L}_{k}(X_{0},\dots,X_{M})\right\|^{2}+\delta_{\widetilde{\Gamma}_{+}^{p}}(X_{0})+\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{1})+\dots+\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{M}),$
@@
$\min_{X=(X_{0}\dots,X_{M})}\leavevmode\nobreak\ \Theta(X_{0},X_{1},\dots,X_{M}).$ (32)
@@
$\partial_{X_{i}}\Theta(X_{0},X_{1},\dots,X_{M})$ $=-T_{2}(V\otimes V){\rm vec}(Q)-T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}\Omega_{k}$
$+2T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}(-\Psi)\mathcal{L}_{k}(X_{0},\dots,X_{i-1},X_{i+1},\dots,X_{M})$
$+2T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}(-\Psi)D_{1}^{\top}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]^{\top}T_{2}^{\top}X_{i}$
$+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{i}).$
@@
$H_{i}(X_{0},\dots,X_{i-1},X_{i+1},\dots,X_{M})=T_{2}(V\otimes V){\rm vec}(Q)+T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}\Omega_{k}$
$-2T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}(-\Psi)\mathcal{L}_{k}(X_{0},\dots,X_{i-1},X_{i+1},\dots,X_{M}),$
$L_{i}=2T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}(-\Psi)D_{1}^{\top}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]^{\top}T_{2}^{\top},$
@@
$\partial_{X_{i}}\Theta(X_{0},X_{1},\dots,X_{M})=-H_{i}(X_{0},\dots,X_{i-1},X_{i+1},\dots,X_{M})+L_{i}X_{i}+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{i}).$
@@
$\bar{X}_{i}^{k+1}=\underset{X_{i}}{\operatorname*{argmin}}\leavevmode\nobreak\ \Theta(X_{0}^{k},X_{1}^{k},\dots,X_{i-1}^{k},X_{i},\bar{X}_{i+1}^{k+1},\dots,\bar{X}_{M}^{k+1})+\frac{1}{2}\left\|X_{i}-X_{i}^{k}\right\|_{\mathcal{S}_{i}}^{2},$
@@
$\mathcal{S}_{i}(X)=\rho_{i}\mathcal{I}(X)-L_{i}X,\leavevmode\nobreak\ \rho_{i}=\max\{{\rm eig}(L_{i})\}$
@@
00 0 $\in-H_{i}(X_{0}^{k},X_{1}^{k},\dots,X_{i-1}^{k},\bar{X}_{i+1}^{k+1},\dots,\bar{X}_{M}^{k+1})+L_{i}X_{i}+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{i})+\mathcal{S}_{i}(X_{i}-X_{i}^{k})$
$=\rho_{i}X_{i}+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{i})-H_{i}(X_{0}^{k},X_{1}^{k},\dots,X_{i-1}^{k},\bar{X}_{i+1}^{k+1},\dots,\bar{X}_{M}^{k+1})-\rho_{i}X_{i}^{k}+L_{i}X_{i}^{k}.$
@@
$\Delta_{i}^{k}=\rho_{i}^{-1}(H_{i}(X_{0}^{k},X_{1}^{k},\dots,X_{i-1}^{k},\bar{X}_{i+1}^{k+1},\dots,\bar{X}_{M}^{k+1})+\rho_{i}X_{i}^{k}-L_{i}X_{i}^{k}),$
@@
$L_{0}=-2\Psi,\leavevmode\nobreak\ H_{0}(X_{1},\dots,X_{M})=-2\Psi\mathcal{L}_{k}(X_{1},\dots,X_{M})-\Omega_{k}.$
@@
$X_{0}^{k+1}=\underset{X_{0}}{\operatorname*{argmin}}\leavevmode\nobreak\ \Theta(X_{0},\bar{X}_{1}^{k+1},\dots,\bar{X}_{M}^{k+1})+\frac{1}{2}\left\|X_{0}-X_{0}^{k}\right\|_{\mathcal{S}_{0}}^{2},$
@@
$\mathcal{S}_{0}(X)=\rho_{0}\mathcal{I}(X)-L_{0}X,\leavevmode\nobreak\ \rho_{0}=\max\{{\rm eig}(L_{0})\}.$
@@
$\tilde{\Delta}_{i}^{k}=\alpha_{i}^{-1}(H_{i}(X_{0}^{k+1},X_{1}^{k+1},\dots,X_{i-1}^{k+1},\bar{X}_{i+1}^{k+1},\dots,\bar{X}_{M}^{k+1})+\rho_{i}\bar{X}_{i}^{k+1}-L_{i}\bar{X}_{i}^{k+1}).$
@@
Introducing the proximal term is first proposed in [12], and the convergence analysis has been considered in [22]. By utilizing the proximal term $\|X_{i}-X_{i}^{k}\|_{\mathcal{S}_{i}}^{2}$, it is possible to derive the closed form expression for the optimization variable $\bar{X}_{i}^{k+1}$. However, a fuller understanding of the proximal term can be achieved by examining it from the perspective of auxiliary functions [22].
@@
$\min f(x)=g(x_{1},\dots,x_{K})+\sum_{i=1}^{K}h_{i}(x_{i})\leavevmode\nobreak\ {\rm s.t.}\leavevmode\nobreak\ x_{i}\in\mathcal{X}_{i},\leavevmode\nobreak\ i=1,\dots,K,$
where $g(\cdot)$is a smooth convex function and $h_{i}$ is a closed convex function. We call that $\{u_{i}(\cdot,\cdot)\}$ is strict auxiliary function, if
@@
    $u_{i}(y_{i},y)=g(y)\leavevmode\nobreak\ \leavevmode\nobreak\ \forall y\in\mathcal{X}=\mathcal{X}_{1}\times\dots\times\mathcal{X}_{K},\forall i$,
@@
    $u_{i}(x_{i},y)\geq g(y_{1},\dots,y_{i-1},x_{i},y_{i+1},\dots,y_{n})\leavevmode\nobreak\ \leavevmode\nobreak\ \forall x_{i}\in\mathcal{X}_{i},\forall y\in\mathcal{X},\forall i,$
@@
    $u_{i}(x_{i},y)\geq u_{i}(\hat{x}_{i},y)+\left\langle\nabla u_{i}(\hat{x}_{i},y),x_{i}-\hat{x}_{i}\right\rangle+\frac{\gamma_{i}}{2}\|x_{i}-\hat{x}_{i}\|^{2},\forall x_{i},\hat{x}_{i}\in\mathcal{X}_{i},$
@@
    For any given $y\in\mathcal{X}$, $u_{k}(x_{i},y)$ has Lipschitz continuous gradient, i.e.,
    $\|\nabla u_{i}(x_{i},y)-\nabla u_{i}(\hat{x}_{i},y)\|\leq L_{i}\|x_{k}-\hat{x}_{i}\|,\leavevmode\nobreak\ \forall\hat{x}_{i},x_{i}\in\mathcal{X}_{i},\forall i,$
    where $L_{i}>0$ is some constant. Further, we have
    $\|\nabla u_{i}(x_{i},y)-\nabla u_{i}(x_{i},z)\|\leq G_{i}\|y-z\|,\leavevmode\nobreak\ \forall x_{i}\in\mathcal{X}_{i},\forall i,\forall y,z\in\mathcal{X}.$

### v3

@@ 4 Solving Inner Layer Optimization Problem
$\mathcal{L}({\rm svec}(W);X)$
$=$ $\langle\tilde{d}_{k},{\rm svec}(W)\rangle+\sigma_{1}\left\|\mathcal{A}D_{1}{\rm svec}(W)+\tilde{b}_{k}\right\|^{2}$
$+\sigma_{2}\left\|D_{1}{\rm svec}(W)-\widetilde{v}_{k}\right\|^{2}$
@@
$\theta(X)$ $=\min_{{\rm svec}(W)}\mathcal{L}({\rm svec}(W);X)$ (35)
$=\langle X_{1}+\dots+X_{M},T_{2}(V\otimes V){\rm vec}(Q)\rangle$
$~{}~{}~{}+\min_{{\rm svec}(W)}\hat{\theta}({\rm svec}(W))$
@@
$\hat{\theta}({\rm svec}(W))=$ $\langle\tilde{d}_{k},{\rm svec}(W)\rangle+\sigma_{1}\left\|\mathcal{A}D_{1}{\rm svec}(W)+\tilde{b}_{k}\right\|^{2}$
$+\sigma_{2}\left\|D_{1}{\rm svec}(W)-\widetilde{v}_{k}\right\|^{2}.$
Letting $\nabla\hat{\theta}({\rm svec}(W))=0$, and denoting
$M=2\sigma_{1}(\mathcal{A}D_{1})^{\top}(\mathcal{A}D_{1})+2\sigma_{2}D_{1}^{\top}D_{1},$
$q_{k}=-2\sigma_{1}(\mathcal{A}D_{1})^{\top}\tilde{b}_{k}+2\sigma_{2}D_{1}^{\top}\widetilde{v}_{k},$
@@
Substituting (36) into (35), it holds
$~{}~{}~{}~{}\theta(X)$
$=\langle X_{1}+\dots+X_{M},T_{2}(V\otimes V){\rm vec}(Q)\rangle$
$~{}~{}~{}+\left\langle\tilde{d}_{k},M^{-1}(q_{k}-\tilde{d}_{k})\right\rangle$
$~{}~{}~{}+\sigma_{1}\left\|\mathcal{A}D_{1}M^{-1}(q_{k}-\tilde{d}_{k})+\tilde{b_{k}}\right\|^{2}$
$~{}~{}~{}+\sigma_{2}\left\|D_{1}M^{-1}(q_{k}-\tilde{d}_{k})-\widetilde{v}_{k}\right\|^{2}$
$=\langle X_{1}+\dots+X_{M},T_{2}(V\otimes V){\rm vec}(Q)\rangle$
$~{}~{}~{}+\tilde{d}_{k}^{\top}\left(\sigma_{1}(\mathcal{A}D_{1}M^{-1})^{\top}(\mathcal{A}D_{1}M^{-1})+\sigma_{2}M^{-2}\right.$
$~{}~{}~{}\left.-M^{-1}\right)\tilde{d}_{k}+\left\langle\tilde{d}_{k},M^{-1}q_{k}+2\sigma_{1}(\mathcal{A}D_{1}M^{-1})^{\top}\xi_{k}^{1}\right.$
$~{}~{}~{}\left.+2\sigma_{2}(D_{1}M^{-1})^{\top}\xi_{k}^{2}\right\rangle+\|\xi_{k}^{1}\|^{2}+\|\xi_{k}^{2}\|^{2},$
@@
$\xi_{k}^{1}=-\mathcal{A}D_{1}M^{-1}q_{k}-\tilde{b}_{k},~{}\xi_{k}^{2}=\widetilde{v}_{k}-D_{1}M^{-1}q_{k}.$
@@
$\max_{X=(X_{0},\dots,X_{M})}{}$ $\langle X_{1}+\dots+X_{M},T_{2}(V\otimes V){\rm vec}(Q)\rangle$ (37)
$+\left\langle\tilde{d}_{k},\Omega_{k}\right\rangle+\tilde{d}_{k}^{\top}\Psi\tilde{d}_{k}$
${\rm s.t.}{}$ $X_{0}\in\widetilde{\Gamma}_{+}^{p},$
$X_{i}\in\widetilde{\Gamma}_{+}^{n},~{}i=1,\dots,M.$
@@
$~{}~{}~{}~{}\Theta(X_{0},X_{1},\dots,X_{M})=$
$-\langle X_{1}+\dots+X_{M},T_{2}(V\otimes V){\rm vec}(Q)\rangle$
$-\left\langle\mathcal{L}_{k}(X_{0},\dots,X_{M}),\Omega_{k}\right\rangle$
$+\left\|\sqrt{-\Psi}\mathcal{L}_{k}(X_{0},\dots,X_{M})\right\|^{2}$
$+\delta_{\widetilde{\Gamma}_{+}^{p}}(X_{0})+\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{1})+\dots+\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{M}),$
@@
$\min_{X=(X_{0}\dots,X_{M})}~{}\Theta(X_{0},X_{1},\dots,X_{M}).$ (38)
@@
$~{}~{}~{}\partial_{X_{i}}\Theta(X_{0},X_{1},\dots,X_{M})=$
$-T_{2}(V\otimes V){\rm vec}(Q)-T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}\Omega_{k}$
$+2T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}(-\Psi)$
$~{}~{}~{}~{}\mathcal{L}_{k}(X_{0},\dots,X_{i-1},X_{i+1},\dots,X_{M})$
$+2T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}(-\Psi)$
$~{}~{}~{}~{}D_{1}^{\top}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]^{\top}T_{2}^{\top}X_{i}+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{i}).$
@@
$~{}~{}~{}H_{i}(X_{0},\dots,X_{i-1},X_{i+1},\dots,X_{M})=$
$T_{2}(V\otimes V){\rm vec}(Q)+T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}\Omega_{k}$
$-2T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}(-\Psi)$
$~{}~{}~{}\mathcal{L}_{k}(X_{0},\dots,X_{i-1},X_{i+1},\dots,X_{M}),$
$L_{i}=2T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}(-\Psi)D_{1}^{\top}$
$~{}~{}~{}~{}~{}~{}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]^{\top}T_{2}^{\top},$
@@
$~{}~{}~{}\partial_{X_{i}}\Theta(X_{0},X_{1},\dots,X_{M})$
$=-H_{i}(X_{0},\dots,X_{i-1},X_{i+1},\dots,X_{M})$
$~{}~{}~{}+L_{i}X_{i}+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{i}).$
@@
$\bar{X}_{i}^{k+1}=\underset{X_{i}}{\operatorname*{argmin}}{}$ $\Theta(X_{0}^{k},X_{1}^{k},\dots,X_{i-1}^{k},X_{i},\bar{X}_{i+1}^{k+1},\dots,\bar{X}_{M}^{k+1})$ (39)
$+\frac{1}{2}\left\|X_{i}-X_{i}^{k}\right\|_{\mathcal{S}_{i}}^{2},$
@@
$\mathcal{S}_{i}(X)=\rho_{i}\mathcal{I}(X)-L_{i}X,~{}\rho_{i}=\max\{{\rm eig}(L_{i})\}$
@@
00 0 $\in-H_{i}(X_{0}^{k},X_{1}^{k},\dots,X_{i-1}^{k},\bar{X}_{i+1}^{k+1},\dots,\bar{X}_{M}^{k+1})$
$~{}~{}~{}+L_{i}X_{i}+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{i})+\mathcal{S}_{i}(X_{i}-X_{i}^{k})$
$=\rho_{i}X_{i}+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{i})$
$~{}~{}~{}-H_{i}(X_{0}^{k},X_{1}^{k},\dots,X_{i-1}^{k},\bar{X}_{i+1}^{k+1},\dots,\bar{X}_{M}^{k+1})$
$~{}~{}~{}-\rho_{i}X_{i}^{k}+L_{i}X_{i}^{k}.$
@@
$\Delta_{i}^{k}=$ $\rho_{i}^{-1}(H_{i}(X_{0}^{k},X_{1}^{k},\dots,X_{i-1}^{k},\bar{X}_{i+1}^{k+1},\dots,\bar{X}_{M}^{k+1})$
$+\rho_{i}X_{i}^{k}-L_{i}X_{i}^{k}),$
@@
Remark 11.
Due to the proximal term $\left\|X_{i}-X_{i}^{k}\right\|_{\mathcal{S}_{i}}^{2}$, optimization problem (39) admits a closed-form solution (40). In fact, if the proximal term is removed, i.e., the update of $\bar{X}_{i}^{k+1}$ is given by
$\bar{X}_{i}^{k+1}=\underset{X_{i}}{\operatorname*{argmin}}~{}\Theta(X_{0}^{k},\dots,X_{i},\dots,\bar{X}_{M}^{k+1}),$ (41)
then Algorithm 2 given below still converges to the global minimizer of (38). However, optimization problem (41) has no closed-form solution, necessitating the development of an iterative algorithm for solving problem (41), e.g., gradient descent (GD) or Nesterov GD. It is worth to highlight that the extra iterative algorithm for problem (41) will significantly improve the complexity of Algorithm 2, and hence the introduction of special proximal term $\left\|X_{i}-X_{i}^{k}\right\|_{\mathcal{S}_{i}}^{2}$ is necessary.
@@
$L_{0}=-2\Psi,~{}H_{0}(X_{1},\dots,X_{M})=-2\Psi\mathcal{L}_{k}(X_{1},\dots,X_{M})-\Omega_{k}.$
@@
$X_{0}^{k+1}=\underset{X_{0}}{\operatorname*{argmin}}~{}\Theta(X_{0},\bar{X}_{1}^{k+1},\dots,\bar{X}_{M}^{k+1})+\frac{1}{2}\left\|X_{0}-X_{0}^{k}\right\|_{\mathcal{S}_{0}}^{2},$
@@
$\mathcal{S}_{0}(X)=\rho_{0}\mathcal{I}(X)-L_{0}X,~{}\rho_{0}=\max\{{\rm eig}(L_{0})\}.$
@@
$\tilde{\Delta}_{i}^{k}=$ $\alpha_{i}^{-1}(H_{i}(X_{0}^{k+1},X_{1}^{k+1},\dots,X_{i-1}^{k+1},\bar{X}_{i+1}^{k+1},\dots,\bar{X}_{M}^{k+1})$
$+\rho_{i}\bar{X}_{i}^{k+1}-L_{i}\bar{X}_{i}^{k+1}).$
@@
Introducing the proximal term for solving separable optimization problem is first proposed in [51], and the convergence analysis can be considered by the analogous methodology of [52], which originally studies the proximal terms in variational inequalities. Remarkably, references [51, 52] only show the fact that introducing the proximal terms has no influence on the algorithm convergence, while how to design the suitable proximal terms remains unstudied, necessitating a case-by-case analysis of individual optimization problem. Hence, the proximal term $\|X_{i}-X_{i}^{k}\|_{\mathcal{S}_{i}}^{2}$ in this paper does not exhibit parallel generalizations of those of [51, 52], but requires additional efforts with consideration of the special structure of dual function $\Theta$; see Remark 11 for details. Additionally, a full understanding of the proximal term can be achieved by examining it from the perspective of auxiliary functions and BSUM framework [53].
Definition 2.
  1. 1.
    $u_{i}(y_{i},y)=f(y)~{}~{}\forall y\in\mathcal{X},\forall i$,
  2. 2.
    $u_{i}(x_{i},y)\geq f(y_{1},\dots,y_{i-1},x_{i},y_{i+1},\dots,y_{n})$ $\forall x_{i}\in\mathcal{X}_{i}$, $\forall y\in\mathcal{X},\forall i,$
  3. 3.
    $u_{i}^{\prime}(x_{i},y;d_{i})|_{x_{i}=y_{i}}=f^{\prime}(y;d)$ $\forall d=(0,\dots,d_{i},\dots,0)$, s.t. $y_{i}+d_{i}\in\mathcal{X}_{i}$ $\forall i$,
  4. 4.
    $u_{i}(x_{i},y)$ is continuous in $(x_{i},y)$ $\forall i$.
Remark 13.
The following separable (possibly nonconvex) problem is considered in [53]:
$\min~{}f(x),~{}~{}{\rm s.t.~{}}x\in\mathcal{X},$ (44)
where $x=(x_{1},\dots,x_{n})^{\top}\in\mathbb{R}^{m}$, and $\mathcal{X}\subseteq\mathbb{R}^{m}$ is the Cartesian product of $n$ closed convex sets: $\mathcal{X}=\mathcal{X}_{1}\times\dots\times\mathcal{X}_{n}$ with $\mathcal{X}_{i}\in\mathbb{R}^{m_{i}}$ and $\sum_{i}m_{i}=m$. The following pseudocode of BSUM framework proposed in [53] gives methodology for solving separable optimization problem (44), where $\{u_{i}\}_{i=1}^{n}$ is a sequence of auxiliary functions of $f$.
Algorithm 1 BSUM
  Given : Feasible initial point $x^{0}\in\mathcal{X}$
  Result: $x^{r}$
  Set $r=0$
  for $k=0,1,2,...$ do
     Let $r=r+1,i=(r\mod n)+1$
     Let $\mathcal{X}^{r}=\underset{x_{i}\in\mathcal{X}_{i}}{\operatorname*{argmin}}~{}u_{i}(x_{i},x^{r-1})$
     Set $x^{r}_{i}$ to be an arbitrary element in $\mathcal{X}^{r}$
     Set $x_{k}^{r}=x_{k}^{r-1}~{}\forall k\neq i$
     if Stopping Criterion==absent$=$ =True then
        return  $x^{r}$
     end if
  end for
We use the notations
$h_{0}(X_{0})=\langle X_{0},\Omega_{k}\rangle,$
$h_{i}(X_{i})=-\left\langle X_{i},T_{2}(V^{\top}\otimes V){\rm vec}(Q)\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.+T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]T_{1}^{\top}\Omega_{k}\right\rangle,$
for $i=1,\dots,M$. Then, optimization problem (38) is equivalent to the following standard form:
$\min\limits_{X_{0},\dots,X_{M}}$ $\widehat{\Theta}(X_{0},\dots,X_{M})=g(X_{0},\dots,X_{M})+\sum_{i=0}^{M}h_{i}(X_{i})$
${\rm s.t.}{}{}{}$ $X_{0}\times X_{1}\dots\times X_{M}\in\widetilde{\Gamma}_{+}^{p}\times\widetilde{\Gamma}_{+}^{n}\times\dots\times\widetilde{\Gamma}_{+}^{n}.$
Hence, ignoring the forward sGS sweep (i.e., regarding Algorithm 2 as BSUM algorithm with Gauss-Seidel (G-S) rule), let
$u_{i}(X_{i},X^{k})=g(X_{0}^{k},\dots,X_{i},\dots,X_{M}^{k})+\frac{1}{2}\|X_{i}-X_{i}^{k}\|_{\mathcal{S}_{i}}^{2},$
with $X^{k}=(X_{0}^{k},\dots,X_{M}^{k})$, for $i=0,\dots,M$, which leads us to derive the following corollary.
@@
$\min{}{}$ $f(x)=g(x_{1},\dots,x_{K})+\sum_{i=1}^{K}h_{i}(x_{i})$
${\rm s.t.}{}{}$ $x_{i}\in\mathcal{X}_{i},~{}i=1,\dots,K,$
where $g$is a smooth convex function and $h_{i}$ is a closed convex function. We call that $\{u_{i}\}$ is a sequence of strict auxiliary functions, if
@@
    $u_{i}(y_{i},y)=g(y)~{}~{}\forall y\in\mathcal{X}=\mathcal{X}_{1}\times\dots\times\mathcal{X}_{K},\forall i$,
@@
    $u_{i}(x_{i},y)\geq g(y_{1},\dots,y_{i-1},x_{i},y_{i+1},\dots,y_{n})~{}~{}\forall x_{i}\in\mathcal{X}_{i},\forall y\in\mathcal{X},\forall i,$
@@
    $u_{i}(x_{i},y)\geq$ $u_{i}(\hat{x}_{i},y)+\left\langle\nabla u_{i}(\hat{x}_{i},y),x_{i}-\hat{x}_{i}\right\rangle$
    $+\frac{\gamma_{i}}{2}\|x_{i}-\hat{x}_{i}\|^{2},\forall x_{i},\hat{x}_{i}\in\mathcal{X}_{i},$
@@
    For any given $y\in\mathcal{X}$, $u_{k}(x_{i},y)$ has Lipschitz continuous gradient, i.e., $\forall\hat{x}_{i},x_{i}\in\mathcal{X}_{i},\forall i,$
    $\|\nabla u_{i}(x_{i},y)-\nabla u_{i}(\hat{x}_{i},y)\|\leq L_{i}\|x_{k}-\hat{x}_{i}\|,$
    where $L_{i}>0$ is some constant. Further, we have $\forall i,\forall y,z\in\mathcal{X}$
    $\|\nabla u_{i}(x_{i},y)-\nabla u_{i}(x_{i},z)\|\leq G_{i}\|y-z\|,~{}\forall x_{i}\in\mathcal{X}_{i}.$